OpenStudy (anonymous):

(6x^-2)^2(0.5x)^4 I dont want a direct answer, please help explain it to me.

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

will medal

OpenStudy (anonymous):

I have tried to solve it and so far I have: 36x^-4

OpenStudy (mathstudent55):

\(\large (6x^{-2})^2(0.5x)^4\) You need to know the rules of exponents. Here are a few you need for this problem. \((ab)^m = a^mb^m\) \((a^m)^n = a^{mn} \)

OpenStudy (mathstudent55):

Start with the left parentheses. You have \(6x^{-2} \) raised to the second power. Use the first rule. To raise a product to a power, raise each factor of the product to the power.

OpenStudy (anonymous):

so 6^2(x^-2)^2

OpenStudy (mathstudent55):

Can you work on the left parentheses and show me what you get? 6x^(-2) is the product of 6 and x^(-2) That product is being raised to the 2nd power. Raise each factor to the 2nd power. Raise 6 to the second power and raise x^-2 to the second power.

OpenStudy (mathstudent55):

Oh, I see. Correct.

OpenStudy (anonymous):

whew glad i got something right today. lol

OpenStudy (mathstudent55):

Now look at the second set of parentheses. Again, you have a product raised to a power. Raise each factor to the power.

OpenStudy (anonymous):

so : (0.5^4x^4)^4 ?

OpenStudy (mathstudent55):

Since you are raising each factor to the power, you don't raise the entire thing to the power again. It's just (0.5)^4(x^4)

OpenStudy (anonymous):

Oh! ok

OpenStudy (mathstudent55):